# WAFT: Warping-Alone Field Transforms for Optical Flow

### v2

@@ Abstract
We introduce Warping-Alone Field Transforms (WAFT), a simple and effective method for optical flow. WAFT is similar to RAFT but replaces cost volume with high-resolution warping, achieving better accuracy with lower memory cost. This design challenges the conventional wisdom that constructing cost volumes is necessary for strong performance. WAFT is a simple and flexible meta-architecture with minimal inductive biases and reliance on custom designs. Compared with existing methods, WAFT ranks 1st on Spring, Sintel, and KITTI benchmarks, achieves the best zero-shot generalization on KITTI, while being up to $4.1\times$ faster than methods with similar performance. Code and model weights are available at https://github.com/princeton-vl/WAFT.
@@ 1 Introduction
Optical flow is a fundamental level vision task that estimates per-pixel 2D motion between video frames. It has many downstream applications, including 3D reconstruction and synthesis (Ma et al., 2022; Zuo & Deng, 2022), action recognition (Sun et al., 2018b; Piergiovanni & Ryoo, 2019; Zhao et al., 2020b), frame interpolation (Xu et al., 2019; Liu et al., 2020; Huang et al., 2020), and autonomous driving (Geiger et al., 2013; Menze & Geiger, 2015; Janai et al., 2020).
Cost volumes (Sun et al., 2018a; Ilg et al., 2017) with iterative updates (Teed & Deng, 2020; Wang et al., 2024) has become a standard design in most state-of-the-art methods (Sun et al., 2018a; Dosovitskiy et al., 2015; Xu et al., 2017; Teed & Deng, 2020; Huang et al., 2022; Wang et al., 2024; Morimitsu et al., 2025), especially when both accuracy and efficiency are taken into account. Previous work (Sun et al., 2018a; Teed & Deng, 2020) regards cost volumes as a more effective representation than image features, as it explicitly models the visual similarity between pixels.
@@
[…]
@@ 2 Related Work
Traditional methods treated optical flow as a global optimization problem that maximizes visual similarity between corresponding pixels (Horn & Schunck, 1981; Zach et al., 2007; Chen & Koltun, 2016; Brox et al., 2004). These methods apply coarse-to-fine warping (Brox et al., 2004; Black & Anandan, 1996; Memin & Perez, 1998), a strategy theoretically justified by Brox et al(Brox et al., 2004), to solve this optimization.
@@
Today, this field is dominated by deep learning methods (Ilg et al., 2017; Dosovitskiy et al., 2015; Sun et al., 2018a; Zhao et al., 2020a; Hui et al., 2018; Teed & Deng, 2020; Sui et al., 2022; Sun et al., 2022; Deng et al., 2023; Huang et al., 2022; Shi et al., 2023; Weinzaepfel et al., 2022; 2023; Xu et al., 2022; 2023b; Leroy et al., 2023; Saxena et al., 2024; Jahedi et al., 2024; 2023; Luo et al., 2022; Zheng et al., 2022; Zhao et al., 2022; Luo et al., 2023; Jung et al., 2023; Luo et al., 2024; Zhou et al., 2024; Morimitsu et al., 2025), which can be categorized into two paradigms: direct or iterative. Direct methods (Dosovitskiy et al., 2015; Weinzaepfel et al., 2022; 2023; Saxena et al., 2024; Leroy et al., 2023; Xu et al., 2022) treat flow estimation as a standard dense prediction task (e.g. monocular depth estimation) and directly regress the dense flow field from large-scale pre-trained models. Iterative methods (Teed & Deng, 2020; Wang et al., 2024; Luo et al., 2024; Morimitsu et al., 2025; Zhou et al., 2024; Sun et al., 2018a; Huang et al., 2022) align more closely with traditional warping-based approaches, refining the flow predictions progressively. Most state-of-the-art methods (Teed & Deng, 2020; Wang et al., 2024; Huang et al., 2022; Shi et al., 2023; Luo et al., 2024; Morimitsu et al., 2025) follow the iterative paradigm due to its significantly higher efficiency than the direct ones.
@@
Cost volumes (Sun et al., 2018a; Dosovitskiy et al., 2015) have been regarded as a standard design in iterative methods (Teed & Deng, 2020; Wang et al., 2024; Morimitsu et al., 2025; Huang et al., 2022; Shi et al., 2023; Xu et al., 2023b; Luo et al., 2024). Prior work (Teed & Deng, 2020; Sun et al., 2018a) empirically shows the effectiveness of cost volumes in handling large displacements. Many iterative methods (Teed & Deng, 2020; Wang et al., 2024; Morimitsu et al., 2025) adopt partial cost volumes (Sun et al., 2018a) to avoid the quadratic computational complexity of full 4D cost volumes; they restrict the search range of each pixel in frame 1 to the neighborhood of its corresponding pixel in frame 2. However, they still suffer from the high memory consumption inherent in cost volumes (Xu et al., 2023a; Zhao et al., 2024).
@@
WAFT is a warping-based iterative method. We achieve state-of-the-art performance across various benchmarks without constructing cost volumes, challenging the conventional wisdom established by previous work (Sun et al., 2018a; Teed & Deng, 2020; Huang et al., 2022). Warping no longer suffers from high memory consumption inherent in cost volumes, which enables high-resolution indexing and therefore improves accuracy.
@@ 3.1 Iterative Methods with Cost Volume
Given two adjacent RGB frames, optical flow predicts pixel-wise 2D motion between adjacent frames. Current iterative methods (Teed & Deng, 2020; Huang et al., 2022; Wang et al., 2024) consist of two parts: (1) input encoders that extract dense image features at low resolution, and (2) a recurrent update module that iteratively refines the flow estimate.
@@
where the operator “concat” concatenates all values inside the set into a vector. In practice, partial cost volumes are usually constructed at multiple scales (Teed & Deng, 2020; Wang et al., 2024) to improve the prediction of large displacements. Current methods also introduce context encoder (Sun et al., 2018a; Teed & Deng, 2020; Huang et al., 2022; Wang et al., 2024) to enhance the effectiveness of iterative refinement.
@@ 3.2 Drawbacks of Cost Volumes
Warping was widely used in both classical and early deep learning approaches (Memin & Perez, 1998; Brox et al., 2004; Ilg et al., 2017; Ranjan & Black, 2017). However, most recent iterative methods (Wang et al., 2024; Teed & Deng, 2020; Huang et al., 2022) have either replaced warping with cost volumes or used both in combination, since cost volumes have been shown to remarkably improve the performance (Sun et al., 2018a). In this section, we analyze the similarities and differences between warping and cost volume (see Figure 2), and argue that with appropriate designs, warping-based methods can achieve performance on par with cost-volume-based methods.
@@ 4 Method
Similar to existing iterative methods (Teed & Deng, 2020; Wang et al., 2024; Huang et al., 2022), our recurrent update module $R$ iteratively predicts the residual flow updates. At step $t$, we concatenate $F(I_{1})$ (feature of frame 1), $\texttt{Warp}(f_{\text{cur}})$ (warped feature of frame 2, see Section 3.3), and the current hidden state $\text{Hidden}_{t}\in\mathbb{R}^{h\times w\times d}$ as input. We use a slightly modified DPT (Ranftl et al., 2021) as the architecture of the module.
@@
We adopt the Mixture-of-Laplace (MoL) loss used in SEA-RAFT (Wang et al., 2024). At step $t$, the hidden state $\text{Hidden}_{t}$ is used to predict the MoL parameters $M\in\mathbb{R}^{h\times w\times 6}$. They are upsampled to the original image resolution through convex upsampling (Teed & Deng, 2020; Wang et al., 2024).
@@
We replace cost volumes, which are standard in existing iterative methods (Sun et al., 2018a; Teed & Deng, 2020; Huang et al., 2022; Wang et al., 2024), with high-resolution warping, which is more memory-efficient. In addition, we have removed the context encoder (Sun et al., 2018a), another flow-specific design standard in existing iterative methods.
@@ 5 Experiments
We report results on Sintel (Butler et al., 2012), KITTI (Geiger et al., 2013), and Spring (Mehl et al., 2023). Following existing work (Teed & Deng, 2020; Huang et al., 2022; Wang et al., 2024; Morimitsu et al., 2025), for training, we use FlythingChairs (Dosovitskiy et al., 2015), FlyingThings (Mayer et al., 2016), HD1K (Kondermann et al., 2016), Sintel (Butler et al., 2012), KITTI (Geiger et al., 2013), Spring (Mehl et al., 2023), and TartanAir (Wang et al., 2020). We use the widely adopted metrics: endpoint-error (EPE), 1-pixel outlier rate (1px), percentage of flow outliers (Fl), and weighted area under the curve (WAUC). Definitions can be found in (Richter et al., 2017; Mehl et al., 2023; Geiger et al., 2013; Morimitsu et al., 2025).
@@ 5.1 Architecture Details
RAFT (Teed & Deng, 2020) 1.61 2.86 5.10 3.07 938 141
@@
Iterative w/ Warping SpyNet (Ranjan & Black, 2017) 6.64 8.36 35.07 26.71 167 25
@@
SpyNet (Ranjan & Black, 2017) 29.963 4.162 12.866 67.150
@@
RAFT (Teed & Deng, 2020) 6.790 1.476 3.198 90.920
@@ 5.3 Benchmark Results
Results are shown in Table 2. Using a Twins backbone only pre-trained on ImageNet, WAFT ranks second on KITTI and is competitive on Sintel (clean). It outperforms prior cost-volume-based SOTA Flowformer++ (Shi et al., 2023) in both accuracy and efficiency given the same backbone, demonstrating the strength of high-resolution warping. The performance can be further improved with stronger backbones (Yang et al., 2024; Siméoni et al., 2025). Using a depth-pretrained DAv2 (Yang et al., 2024), on KITTI (Geiger et al., 2013), WAFT achieves the best Fl on non-occluded pixels and the second best on all pixels. It also ranks first on Sintel (Clean) (Butler et al., 2012) and competitive on Sintel (Final). WAFT is up to $4.1\times$ faster than comparable methods, demonstrating its high efficiency.
@@
It appears that warping as a network operation has been largely abandoned by works in the last 8 years and the last time warping-based methods achieved top positions on the leaderboards were around 2017 (Ranjan & Black, 2017; Ilg et al., 2017). WAFT is significant in that it has revisited and revived an idea that has fallen out of favor. Compared to methods that do use warping (Ranjan & Black, 2017; Ilg et al., 2017), WAFT reduces endpoint-error (EPE) by at least 64% on Sintel and 70% on Spring (Mehl et al., 2023), while also reducing Fl by at least 68% on KITTI (Geiger et al., 2013) and 57% on Spring (Mehl et al., 2023). Besides, WAFT reduces 1px-outlier rate by 52% on Spring (Mehl et al., 2023).
@@ 5.4 Zero-Shot Evaluation
RAFT (Teed & Deng, 2020) 1.43 2.71 5.04 17.4
@@
Following previous work (Teed & Deng, 2020; Huang et al., 2022; Sun et al., 2018a), we train our model on FlyingChairs (Dosovitskiy et al., 2015) and FlyingThings (Mayer et al., 2016). Then we evaluate the performance on the training split of Sintel (Butler et al., 2012) and KITTI (Geiger et al., 2013).
@@ 5.5 Ablation Study
The vision transformer design is crucial to iterative warping. We observe a significant performance drop when replacing the DPT-based recurrent update module with CNNs, highlighting the importance of modeling long-range dependence. This finding may help explain why early deep learning approaches (Ilg et al., 2017; Ranjan & Black, 2017) that implemented warping using CNNs underperformed compared to cost-volume-based methods (Sun et al., 2018a; Teed & Deng, 2020; Huang et al., 2022).
@@
We also design a cost-volume-based variant following the common setup (Wang et al., 2024; Teed & Deng, 2020; Huang et al., 2022) with look-up radius $4$ at 1/8 resolution, and find that it performs similarly to the warping counterpart (shown in Table 5) but costs $2.2\times$ training memory (21.2 GiB vs. 9.5 GiB).
@@
Iterative updates achieve better performance than direct regression within our meta-architecture. We implement direct regression by setting the number of iterations $T=1$. For fair comparison, we scale up the networks to match the computational cost of 5-iteration WAFT. Our results show that WAFT significantly outperforms these direct regression variants, indicating the effectiveness and high efficiency of the iterative paradigm. This finding aligns with the observation that existing direct methods either underperform (Xu et al., 2022; Weinzaepfel et al., 2022) or require substantially more computational cost (Saxena et al., 2024; Weinzaepfel et al., 2022; 2023) compared to iterative approaches (Teed & Deng, 2020; Wang et al., 2024; Huang et al., 2022; Morimitsu et al., 2025).
@@
Feature-space warping is more effective than image-space warping, which is commonly used in classic methods (Ma et al., 2022; Brox et al., 2004; Black & Anandan, 1996) and early deep learning methods (Ilg et al., 2017; Ranjan & Black, 2017). Feature-space warping does not need to re-extract features for the warped image in each iteration, significantly saving computational cost while achieving slightly better accuracy.
@@
Prior work (Sun et al., 2018a; Teed & Deng, 2020; Wang et al., 2024; Huang et al., 2022) has often used a context encoder that provides an extra input to the update module. Our ablation show that the context encoder is not necessary. The context encoder introduces additional computation overhead, but does not significantly affect performance. Previous work (Wang et al., 2024) also points out that the context encoder can be regarded as a direct flow regressor, which functions similarly to the first iteration of WAFT.

### v3

@@ Abstract
We introduce Warping-Alone Field Transforms (WAFT), a simple and effective method for optical flow. WAFT is similar to RAFT but replaces cost volume with high-resolution warping, achieving better accuracy with lower memory cost. This design challenges the conventional wisdom that constructing cost volumes is necessary for strong performance. WAFT is a simple and flexible meta-architecture with minimal inductive biases and reliance on custom designs. Compared with existing methods, WAFT ranks 1st on Spring, Sintel, and KITTI benchmarks, achieves the best zero-shot generalization on KITTI, while being $1.3-4.1\times$ faster than existing methods that have competitive accuracy (e.g., $1.3\times$ than Flowformer++, $4.1\times$ than CCMR+). Code and model weights are available at https://github.com/princeton-vl/WAFT.
@@ 1 Introduction
Optical flow is a fundamental low-level vision task that estimates per-pixel 2D motion between video frames. It has many downstream applications, including 3D reconstruction and synthesis (Ma et al., 2022; Zuo and Deng, 2022), action recognition (Sun et al., 2018b; Piergiovanni and Ryoo, 2019; Zhao et al., 2020b), frame interpolation (Xu et al., 2019; Liu et al., 2020; Huang et al., 2020), and autonomous driving (Geiger et al., 2013; Menze and Geiger, 2015; Janai et al., 2020).
Cost volumes (Sun et al., 2018a; Ilg et al., 2017) with iterative updates (Teed and Deng, 2020; Wang et al., 2024) has become a standard design in most state-of-the-art methods (Sun et al., 2018a; Dosovitskiy et al., 2015; Xu et al., 2017; Teed and Deng, 2020; Huang et al., 2022; Wang et al., 2024; Morimitsu et al., 2025), especially when both accuracy and efficiency are taken into account. Previous work (Sun et al., 2018a; Teed and Deng, 2020) regards cost volumes as a more effective representation than image features, as it explicitly models the visual similarity between pixels.
@@
WAFT achieves state-of-the-art performance across various benchmarks with high efficiency and a simple design. Using a Twins (Chu et al., 2021) backbone pre-trained only on ImageNet, WAFT ranks first on Spring, second on KITTI, and is competitive on Sintel. It also achieves the best zero-shot cross-dataset generalization on KITTI. Using a stronger backbone, depth-pretrained DAv2 (Yang et al., 2024), WAFT outperforms existing methods on all public benchmarks. We achieve this with standard network architectures for the sub-modules  (Dosovitskiy et al., 2020; He et al., 2016; Ranftl et al., 2021), removing custom designs typically needed in prior work, while being $1.3-4.1\times$ faster than existing methods that have competitive accuracy (e.g., $1.3\times$ than Flowformer++ Shi et al. (2023), $4.1\times$ than CCMR+ Jahedi et al. (2024)).
@@ 2 Related Work
Traditional methods treated optical flow as a global optimization problem that maximizes visual similarity between corresponding pixels (Horn and Schunck, 1981; Zach et al., 2007; Chen and Koltun, 2016; Brox et al., 2004). These methods apply coarse-to-fine warping (Brox et al., 2004; Black and Anandan, 1996; Memin and Perez, 1998), a strategy theoretically justified by Brox et al. (Brox et al., 2004), to solve this optimization.
@@
Today, this field is dominated by deep learning methods (Ilg et al., 2017; Dosovitskiy et al., 2015; Sun et al., 2018a; Zhao et al., 2020a; Hui et al., 2018; Teed and Deng, 2020; Sui et al., 2022; Sun et al., 2022; Deng et al., 2023; Huang et al., 2022; Shi et al., 2023; Weinzaepfel et al., 2022; 2023; Xu et al., 2022; 2023b; Leroy et al., 2023; Saxena et al., 2024; Jahedi et al., 2024; 2023; Luo et al., 2022; Zheng et al., 2022; Zhao et al., 2022; Luo et al., 2023; Jung et al., 2023; Luo et al., 2024; Zhou et al., 2024; Morimitsu et al., 2025), which can be categorized into two paradigms: direct or iterative. Direct methods (Dosovitskiy et al., 2015; Weinzaepfel et al., 2022; 2023; Saxena et al., 2024; Leroy et al., 2023; Xu et al., 2022) treat flow estimation as a standard dense prediction task (e.g. monocular depth estimation) and directly regress the dense flow field from large-scale pre-trained models. Iterative methods (Teed and Deng, 2020; Wang et al., 2024; Luo et al., 2024; Morimitsu et al., 2025; Zhou et al., 2024; Sun et al., 2018a; Huang et al., 2022) align more closely with traditional warping-based approaches, refining the flow predictions progressively. Most state-of-the-art methods (Teed and Deng, 2020; Wang et al., 2024; Huang et al., 2022; Shi et al., 2023; Luo et al., 2024; Morimitsu et al., 2025) follow the iterative paradigm due to its significantly higher efficiency than the direct ones.
@@
Cost volumes (Sun et al., 2018a; Dosovitskiy et al., 2015) have been regarded as a standard design in iterative methods (Teed and Deng, 2020; Wang et al., 2024; Morimitsu et al., 2025; Huang et al., 2022; Shi et al., 2023; Xu et al., 2023b; Luo et al., 2024). Prior work (Teed and Deng, 2020; Sun et al., 2018a) empirically shows the effectiveness of cost volumes in handling large displacements. Many iterative methods (Teed and Deng, 2020; Wang et al., 2024; Morimitsu et al., 2025) adopt partial cost volumes (Sun et al., 2018a) to avoid the quadratic computational complexity of full 4D cost volumes; they restrict the search range of each pixel in frame 1 to the neighborhood of its corresponding pixel in frame 2. However, they still suffer from the high memory consumption inherent in cost volumes (Xu et al., 2023a; Zhao et al., 2024).
@@
WAFT is a warping-based iterative method. We achieve state-of-the-art performance across various benchmarks without constructing cost volumes, challenging the conventional wisdom established by previous work (Sun et al., 2018a; Teed and Deng, 2020; Huang et al., 2022). Warping no longer suffers from high memory consumption inherent in cost volumes, which enables high-resolution indexing and therefore improves accuracy.
@@ 3.1 Iterative Methods with Cost Volume
Given two adjacent RGB frames, optical flow predicts pixel-wise 2D motion between adjacent frames. Current iterative methods (Teed and Deng, 2020; Huang et al., 2022; Wang et al., 2024) consist of two parts: (1) input encoders that extract dense image features at low resolution, and (2) a recurrent update module that iteratively refines the flow estimate.
@@
where the operator “concat” concatenates all values inside the set into a vector. In practice, partial cost volumes are usually constructed at multiple scales (Teed and Deng, 2020; Wang et al., 2024) to improve the prediction of large displacements. Current methods also introduce context encoder (Sun et al., 2018a; Teed and Deng, 2020; Huang et al., 2022; Wang et al., 2024) to enhance the effectiveness of iterative refinement.
@@ 3.2 Drawbacks of Cost Volumes
Warping was widely used in both classical and early deep learning approaches (Memin and Perez, 1998; Brox et al., 2004; Ilg et al., 2017; Ranjan and Black, 2017). However, most recent iterative methods (Wang et al., 2024; Teed and Deng, 2020; Huang et al., 2022) have either replaced warping with cost volumes or used both in combination, since cost volumes have been shown to remarkably improve the performance (Sun et al., 2018a). In this section, we analyze the similarities and differences between warping and cost volume (see Figure 2), and argue that with appropriate designs, warping-based methods can achieve performance on par with cost-volume-based methods.
@@ 4 Method
Similar to existing iterative methods (Teed and Deng, 2020; Wang et al., 2024; Huang et al., 2022), our recurrent update module $R$ iteratively predicts the residual flow updates. At step $t$, we concatenate $F(I_{1})$ (feature of frame 1), $\texttt{Warp}(f_{\text{cur}})$ (warped feature of frame 2, see Section 3.3), and the current hidden state $\text{Hidden}_{t}\in\mathbb{R}^{h\times w\times d}$ as input. We use a slightly modified DPT (Ranftl et al., 2021) as the architecture of the module.
@@
We adopt the Mixture-of-Laplace (MoL) loss used in SEA-RAFT (Wang et al., 2024). At step $t$, the hidden state $\text{Hidden}_{t}$ is used to predict the MoL parameters $M\in\mathbb{R}^{h\times w\times 6}$. They are upsampled to the original image resolution through convex upsampling (Teed and Deng, 2020; Wang et al., 2024).
@@
We replace cost volumes, which are standard in existing iterative methods (Sun et al., 2018a; Teed and Deng, 2020; Huang et al., 2022; Wang et al., 2024), with high-resolution warping, which is more memory-efficient. In addition, we have removed the context encoder (Sun et al., 2018a), another flow-specific design standard in existing iterative methods.
@@ 5 Experiments
We report results on Sintel (Butler et al., 2012), KITTI (Geiger et al., 2013), and Spring (Mehl et al., 2023). Following existing work (Teed and Deng, 2020; Huang et al., 2022; Wang et al., 2024; Morimitsu et al., 2025), for training, we use FlythingChairs (Dosovitskiy et al., 2015), FlyingThings (Mayer et al., 2016), HD1K (Kondermann et al., 2016), Sintel (Butler et al., 2012), KITTI (Geiger et al., 2013), Spring (Mehl et al., 2023), and TartanAir (Wang et al., 2020). We use the widely adopted metrics: endpoint-error (EPE), 1-pixel outlier rate (1px), percentage of flow outliers (Fl), and weighted area under the curve (WAUC). Definitions can be found in (Richter et al., 2017; Mehl et al., 2023; Geiger et al., 2013; Morimitsu et al., 2025).
@@ 5.1 Architecture Details
RAFT (Teed and Deng, 2020) 1.61 2.86 5.10 3.07 938 141
@@
Iterative w/ Warping SpyNet (Ranjan and Black, 2017) 6.64 8.36 35.07 26.71 167 25
@@
SpyNet (Ranjan and Black, 2017) 29.963 4.162 12.866 67.150
@@
RAFT (Teed and Deng, 2020) 6.790 1.476 3.198 90.920
@@ 5.3 Benchmark Results
[…]
@@
It appears that warping as a network operation has been largely abandoned by works in the last 8 years and the last time warping-based methods achieved top positions on the leaderboards were around 2017 (Ranjan and Black, 2017; Ilg et al., 2017). WAFT is significant in that it has revisited and revived an idea that has fallen out of favor. Compared to methods that do use warping (Ranjan and Black, 2017; Ilg et al., 2017), WAFT reduces endpoint-error (EPE) by at least 64% on Sintel and 70% on Spring (Mehl et al., 2023), while also reducing Fl by at least 68% on KITTI (Geiger et al., 2013) and 57% on Spring (Mehl et al., 2023). Besides, WAFT reduces 1px-outlier rate by 52% on Spring (Mehl et al., 2023).
@@ 5.4 Zero-Shot Evaluation
RAFT (Teed and Deng, 2020) 1.43 2.71 5.04 17.4
@@
Following previous work (Teed and Deng, 2020; Huang et al., 2022; Sun et al., 2018a), we train our model on FlyingChairs (Dosovitskiy et al., 2015) and FlyingThings (Mayer et al., 2016). Then we evaluate the performance on the training split of Sintel (Butler et al., 2012) and KITTI (Geiger et al., 2013).
@@ 5.5 Ablation Study
The vision transformer design is crucial to iterative warping. We observe a significant performance drop when replacing the DPT-based recurrent update module with CNNs, highlighting the importance of modeling long-range dependence. This finding may help explain why early deep learning approaches (Ilg et al., 2017; Ranjan and Black, 2017) that implemented warping using CNNs underperformed compared to cost-volume-based methods (Sun et al., 2018a; Teed and Deng, 2020; Huang et al., 2022).
@@
We also design a cost-volume-based variant following the common setup (Wang et al., 2024; Teed and Deng, 2020; Huang et al., 2022) with look-up radius $4$ at 1/8 resolution, and find that it performs similarly to the warping counterpart (shown in Table 5) but costs $2.2\times$ training memory (21.2 GiB vs. 9.5 GiB).
@@
Iterative updates achieve better performance than direct regression within our meta-architecture. We implement direct regression by setting the number of iterations $T=1$. For fair comparison, we scale up the networks to match the computational cost of 5-iteration WAFT. Our results show that WAFT significantly outperforms these direct regression variants, indicating the effectiveness and high efficiency of the iterative paradigm. This finding aligns with the observation that existing direct methods either underperform (Xu et al., 2022; Weinzaepfel et al., 2022) or require substantially more computational cost (Saxena et al., 2024; Weinzaepfel et al., 2022; 2023) compared to iterative approaches (Teed and Deng, 2020; Wang et al., 2024; Huang et al., 2022; Morimitsu et al., 2025).
@@
Feature-space warping is more effective than image-space warping, which is commonly used in classic methods (Ma et al., 2022; Brox et al., 2004; Black and Anandan, 1996) and early deep learning methods (Ilg et al., 2017; Ranjan and Black, 2017). Feature-space warping does not need to re-extract features for the warped image in each iteration, significantly saving computational cost while achieving slightly better accuracy.
@@
Prior work (Sun et al., 2018a; Teed and Deng, 2020; Wang et al., 2024; Huang et al., 2022) has often used a context encoder that provides an extra input to the update module. Our ablation show that the context encoder is not necessary. The context encoder introduces additional computation overhead, but does not significantly affect performance. Previous work (Wang et al., 2024) also points out that the context encoder can be regarded as a direct flow regressor, which functions similarly to the first iteration of WAFT.